Nesmith Corporations outstanding bonds have a $1,000 par value, a 6% semiannual coupon, 18 years to maturity, and a 9% YTM.
Answer #1

No of periods (N) = 18 x 2 = 36
YTM for half year 9/2 = 4.5%
Coupon amount (semi annual )= 1000×6%×(0.5) = 30
Par value = 1000

The price of the bond = coupon amount x PVAF(YTM, N) + Par value× PVIF (YTM, N)

=30 x PVAF(4.5%, 36) + 1000× PVIF (4.5%, 36 )
= $735.01

The price of the bond = $735.01

(Note:
PVAF = present value annuity factor.
PVIF = present value interest factor.)